OpenStudy (anonymous):

If (x+5) is a factor of x^3+2x^2-13x+10, find the other factors.

OpenStudy (anonymous):

@Neemo

OpenStudy (anonymous):

u need factor of x^3+2x^2-13x+10

OpenStudy (anonymous):

\[(x+5)(ax^2+bx+c)\] \[(x+5)(x^2+bx+2)\]

OpenStudy (anonymous):

i need factors of \[x ^{3}+2x ^{2}-13x+10\] other than (x+5)

OpenStudy (anonymous):

\[5x^2+bx^2=2x^2\] \[b=-3\] \[(x+5(x^2-3x+2)\]

OpenStudy (anonymous):

\[(x+5)(x^2-3x+2)=(x+5)(x-2)(x-1)\]

OpenStudy (anonymous):

x^3+2x^2-13x+10 x^3+5x^2-3x^2-15x+2x+10 =x^2 (x+5) -3x(x+5) +2(x+5) =(x+5)(x^2-3x+2) =(x+5)(x^2-2x-x+2) =(x+5)(x(x-2)-1(x-2)) =(x+5)(x-2)(x-1)

OpenStudy (anonymous):

@basith ....do you still need help on this?

OpenStudy (anonymous):

nop.......thnx any way....@yummydum :)

OpenStudy (anonymous):

@yummydum